OpenStudy (anonymous):

Find the Vertex:

OpenStudy (anonymous):

\[y=2x^2-4x+8\]

Directrix (directrix):

Do you know the formula for the x-coordinate of the vertex?

OpenStudy (campbell_st):

use x = -b/2a which gives the line of symmetry x = 4/4 x = 1 is the line of symmetry substitute into the equation y = 6 vertex at (1, 6) ... I think

OpenStudy (anonymous):

\[f(x)=a(x-h)^2+k\]

OpenStudy (campbell_st):

if you use your method you get f(x) = 2(x^2 - 2x) + 8 f(x) = 2(x^2 - 2x +1) + 6 f(x) 2(x -1)^2 + 6 same solution just a bit more difficult.. find the line of symmetry and substitute

OpenStudy (anonymous):

You could always evaluate \(f'\left(x\right)=4x-4=0\therefore x=1\) and find the coordinate \(\left(x,f\left(x\right)\right)\) for that point.